A delivery van is carrying a load of ten flat screen televisions and forty five CD players. One third of the load is flat screen television and one sixth is CD players. If the televisions weigh eleven kilograms each, what is the total weight in the van?
A) 330kg b) 660 kg c) 3110kg d) 990kg

Found 2 solutions by mananth, math_tutor2020:
Answer by mananth(16946)


A delivery van is carrying a load of ten flat screen televisions and forty five CD players. One third of the load is flat screen television and one sixth is CD players. If the televisions weigh eleven kilograms each, what is the total weight in the van?


110=1/3 x
x=330 total weight

Answer by math_tutor2020(3816)


Answer: 330 kg (choice A)


Work Shown:


weight of 1 tv = 11 kg
weight of 10 tvs = 10*11 = 110 kg

w = total weight
1/3 of the total weight is the tvs
1/3 of w = 110 kg
(1/3)w = 110
w/3 = 110
w = 3*110
w = 330 kg is the total weight

It appears the info about the cd players is a distraction.
We don't use this info at all.